# OpenVLThinker: Complex Vision-Language Reasoning via Iterative SFT-RL Cycles

### v2

@@ 5.2 Analysis
At SFT-Iter1, we utilized the base model Qwen2.5-VL-7B to generate image descriptions and obtained R1-like reasoning from QwQ-32B through rejection sampling. A common problem for distillation observed in text-only math reasoning is the overly long reasoning length coupled with unnecessary repetitions of reflections [77, 41]. We observed similarly that these initial reasoning traces were often excessively verbose, partly due to information loss during image-to-caption conversion. Consequently, post-SFT reasoning became increasingly repetitive with unproductive self-reflections (see Appendix C for an illustration). To address this, we evaluated two filtering strategies: (1) discarding samples with reasoning traces exceeding 500 words, and (2) truncating reflections by splitting traces of at specific keywords that were overly repetitive in data ( “Wait,” “But wait,” and “But the question”) and discarding subsequent segments while preserving the final answer. The latter approach was ultimately adopted to prevent the model from internalizing reflection loops, while preserving the reasoning action at a reasonable frequency. Table 5.2 compares models trained on original versus processed data.
@@
We conducted a quantitative analysis to categorize the data sources based on difficulty. Applying k-means clustering (with $k=3$) to our composite difficulty score as described in Section 4 allowed us to clearly identify three distinct difficulty pools, as shown in Table 4. We visualize the difficulty scores for each source in Figure 7. In Figure 8, we show the performance of GRPO-Iter1 when drawing 3K data from either (1) 10 data sources classified as either Easy or Medium, or (2) 5 data sources classified as Medium. We observe that RL training with easy-level data results in ineffective performance gain as compared to sourcing from medium-level data only. This finding aligns with concurrent algorithmic efforts such as DAPO [78] in the text-only domain for improving GRPO by dynamically filtering out overly-easy examples.
@@
Building upon the performance improvements shown in Figure 4, we further analyze changes in reasoning length across iterations, as illustrated in Figure 10. Our results indicate that the reasoning model consistently utilizes more words at inference time compared to the base non-reasoning model, without becoming excessively repetitive. Notably, the largest increase in reasoning length occurs at Iteration 1, with subsequent iterations gradually adopting more concise reasoning. This progression suggests an increasingly efficient utilization of reflective reasoning, engaging reflections primarily when beneficial. In Appendix C (Figure 13 and 14), we show reasoning examples that our SFT-ed model was incorrect while our RL-ed model was correct.
@@ 6 Conclusion
Appendix A Additional Empirical Study
@@
Does Complex Reasoning Matter for VQA? We additionally investigated whether complex, multi-step reasoning provides significant performance gains over standard (non-R1) reasoning in visual tasks. In this study, we use the ConTextual [65] validation set of 100 VQA examples, aiming to disentangle the roles of image grounding and textual reasoning. As similar to our first distillation process, we separately employ a vision-language model for caption generation and a pure-text model for reasoning. The image description generated by the captioning model is then fed into one of two text-based models: DeepSeek-R1-Distill-14B (an R1-style reasoner) or Qwen2.5-14B-Instruct (a standard instruction-tuned model). This setup allows us to isolate the impact of R1 reasoning from the effects of the underlying vision encoder.
@@
Appendix B Experiment Details
We thank LLaMA-Factory [^6] https://github.com/hiyouga/LLaMA-Factory and EasyR1777https://github.com/hiyouga/EasyR1 for open-sourcing the training framework that we used for SFT and GRPO. In Table 7 and 8, we detail the hyperparameters that we used for SFT, GRPO and inference. We further lay out the prompts we used for generating image captions. Experiments were conducted on GPU clusters to the similar level of NVIDIA H100 80GB GPU. SFT/Distillation requires 30 minutes and RL requires 20 hours for each iteration. In addition, distillation data generation with verification requires about 8 hours.
@@ Appendix B Experiment Details
Appendix C Output Examples
@@
Example of SFT-Iter3 incorrect output and GRPO-Iter3 Correct output to the same question Question: The derivative of f(x) at x=2 is ____ that at x=5 Choices: (A) larger than (B) equal to (C) smaller than SFT-Iter3
@@
Okay, let me see. The question is asking about the derivative of the function f(x) = |2x - 3| + 1 at x=2 compared to x=5. First, I need to recall what the derivative represents. The derivative gives the slope of the tangent line at a point on the function. So I need to look at the slopes at those two points. </think> The function ( f(x) = |2x - 3| + 1 ) is an absolute value function, which means it has a V-shape. The vertex of this V-shape occurs where the expression inside the absolute value is zero, i.e., ( 2x - 3 = 0 ), which gives ( x = $\frac{3}{2}$ ). The derivative of ( f(x) ) will be zero at the vertex because the function changes direction there. For ( $x<\frac{3}{2}$ ), the function is decreasing, and for ( $x>\frac{3}{2}$ ), it is increasing. The slopes on either side of the vertex are negative and positive, respectively. At ( x = 2 ) (which is to the right of the vertex), the function is increasing, so the derivative is positive. At ( x = 5 ) (which is further to the right), the function is also increasing, so the derivative is positive but steeper because the function is increasing faster. Since the function is increasing faster at ( x = 5 ) than at ( x = 2 ), the derivative at ( x = 5 ) is larger. <answer>(C)</answer> ================== GRPO-Iter3 <think> Okay, let’s see. The question is asking about the derivative of the function f(x) = |2x - 3| + 1 at x=2 compared to x=5. First, I need to recall what the derivative represents. The derivative at a point gives the slope of the tangent line at that point. So I need to look at the slopes at x=2 and x=5. </think> The function ( f(x) = |2x - 3| + 1 ) is an absolute value function shifted up by 1. The absolute value function ( |2x - 3| ) has a V-shape with a vertex at ( x = $\frac{3}{2}$ ). The derivative of ( |2x - 3| ) is ( 2 ) for ( x $>\frac{3}{2}$ ) and ( -2 ) for ( x < $\frac{3}{2}$ ). Adding 1 to the function doesn’t affect the derivative’s magnitude but shifts the y-values up. At ( x = 2 ), which is greater than ( $\frac{3}{2}$ ), the derivative is ( 2 ). At ( x = 5 ), which is also greater than ( $\frac{3}{2}$ ), the derivative is also ( 2 ). Since the derivative is the same at both points, the slopes are equal. <answer>(B)</answer> Ground truth: (B)
Example of SFT-Iter3 incorrect output and GRPO-Iter3 Correct output to the same question Question: Find $m\angle 3$. Choices: (A) 28 (B) 38 (C) 52 (D) 62 SFT-Iter3
@@
Okay, let me see. The problem is asking for the measure of angle 3 in the given quadrilateral. The hint suggests looking at the angles provided, which are 52 degrees and 38 degrees. Hmm, the quadrilateral is a rectangle, so all angles should be 90 degrees. But the hint is confusing me. Let me check the options again. The angles 52 and 38 are given, but they’re not the ones I need to find. The question is about angle 3, which is opposite angle 1. Since opposite angles in a rectangle are equal, angle 3 should be the same as angle 1. But angle 1 is not given directly. Let me see, the hint is saying to look at the angles provided. The 52 and 38 are supplementary angles, which means they add up to 180 degrees. But angle 3 is not supplementary to 52 or 38. It’s the opposite angle, so it should be the same as angle 1. Since angle 1 is not given, I need to use the options. The only angle that could be the same as angle 1 is the one that’s not given directly. The options are A, B, C, D. The only one that’s not supplementary to 52 or 38 is 62. So angle 3 must be 62 degrees. </think> The problem states that the quadrilateral is a rectangle, and all angles in a rectangle are 90 degrees. The hint provided the angles 52° and 38°, which are supplementary angles (52° + 38° = 90°). However, angle 3 is the opposite angle to angle 1, which means they are equal because opposite angles in a rectangle are congruent. Since angle 1 is not given directly, we need to look at the options provided. The only angle that fits the properties of opposite angles in a rectangle is the one that is not supplementary to the given angles. The options are A (28), B (38), C (52), and D (62). Since angle 1 is not given, the only angle that could be the same as angle 3 is the one that is not supplementary to 52° or 38°. The only angle that fits this criterion is 62°, as it is the only one not listed as supplementary to the given angles. <answer>(D)</answer> ================== GRPO-Iter3 <think> Okay, let’s see. The question is asking for the measure of angle 3. The image provided is a rectangle with some angles labeled. The hint suggests looking at the angles given, which are 52 degrees and 38 degrees. Hmm, the rectangle has right angles, so each angle should be 90 degrees. The angles 52 and 38 are supplementary because they add up to 90 degrees (90 - 52 = 38). But angle 3 is not one of those. It’s the angle opposite the 38-degree angle, so it should be the same. Wait, the question is about angle 3, which is the one opposite the 38-degree angle. So it should be 38 degrees too, right? </think> The problem states that angle 3 is opposite the 38-degree angle. In a rectangle, opposite angles are equal. Since angle 3 is opposite the 38-degree angle, angle 3 must also be 38 degrees. <answer>(B)</answer> Ground truth: (B)

### v3

@@ 5.2 Analysis
At SFT-Iter1, we utilized the base model Qwen2.5-VL-7B to generate image descriptions and obtained R1-like reasoning from QwQ-32B through rejection sampling. A common problem for distillation observed in text-only math reasoning is the overly long reasoning length coupled with unnecessary repetitions of reflections [yeo2025demystifying, liu2025there]. We observed similarly that these initial reasoning traces were often excessively verbose, partly due to information loss during image-to-caption conversion. Consequently, post-SFT reasoning became increasingly repetitive with unproductive self-reflections (see Appendix D for an illustration). To address this, we evaluated two filtering strategies: (1) discarding samples with reasoning traces exceeding 500 words, and (2) truncating reflections by splitting traces of at specific keywords that were overly repetitive in data ( “Wait,” “But wait,” and “But the question”) and discarding subsequent segments while preserving the final answer. The latter approach was ultimately adopted to prevent the model from internalizing reflection loops, while preserving the reasoning action at a reasonable frequency. Table 5.2 compares models trained on original versus processed data.
@@
We conducted a quantitative analysis to categorize the data sources based on difficulty. Applying k-means clustering (with $k=3$) to our composite difficulty score as described in Section 4 allowed us to clearly identify three distinct difficulty pools, as shown in Table 4. We visualize the difficulty scores for each source in Figure 7. In Figure 8, we show the performance of GRPO-Iter1 when drawing 3K data from either (1) 10 data sources classified as either Easy or Medium, or (2) 5 data sources classified as Medium. We observe that RL training with easy-level data results in ineffective performance gain as compared to sourcing from medium-level data only. This finding aligns with concurrent algorithmic efforts such as DAPO [yu2025dapo] in the text-only domain for improving GRPO by dynamically filtering out overly-easy examples.
@@
Building upon the performance improvements shown in Figure 4, we further analyze changes in reasoning length across iterations, as illustrated in Figure 10. Our results indicate that the reasoning model consistently utilizes more words at inference time compared to the base non-reasoning model, without becoming excessively repetitive. Notably, the largest increase in reasoning length occurs at Iteration 1, with subsequent iterations gradually adopting more concise reasoning. This progression suggests an increasingly efficient utilization of reflective reasoning, engaging reflections primarily when beneficial. In Appendix D (Figure 13 and 14), we show reasoning examples that our SFT-ed model was incorrect while our RL-ed model was correct.
Design choice on restarting iterations
Restarting training from scratch at each iteration is a standard practice in iterative self-improvement methods [singh2023beyond, hosseini2024v]. This design choice ensures training stability and prevents overfitting, especially when the data scale is relatively small, thus maintaining better generalization to unseen tasks. As noted in [singh2023beyond], re-training from the base model provides comparable task-specific performance and significantly better transfer to held-out tasks compared to continued training.
In our iterative re-training approach, the SFT data is refined and improved across iterations, while the base model parameters are reinitialized to prevent error accumulation. To further substantiate this design choice, we conducted an additional comparison between (a) re-training from scratch and (b) continuing training from the previous checkpoint.
Table 5: Comparison between re-training from scratch and continued training.
Method MathVista EMMA HallusionBench
OpenVLThinker (re-training) 71.7 25.8 70.2
OpenVLThinker (continue training) 71.8 25.1 66.8
The results indicate that continued training leads to performance degradation on HallusionBench, suggesting potential overfitting to the previous iteration’s data. Hence, restarting from the base model offers a more robust and generalizable learning trajectory across iterations.
@@ 6 Conclusion
Acknowledgments
We thank anonymous reviewers for their helpful comments. This work was partially supported by U.S. DARPA ECOLE Program No. #HR00112390060, ONR grant N00014-23-1-2780, DARPA ANSR program FA8750- 23-2-0004, Amazon, and Apple. Chang and Peng were supported in part by a grant from DARPA to the Simons Institute for the Theory of Computing. Bansal was supported in part by AFOSR MURI grant FA9550-22-1-0380. Wang was supported by National Science Foundation (2106859, 2200274, 2312501), National Institutes of Health (U54HG012517, U24DK097771, U54OD036472), NEC, and Optum AI.
Appendix A Additional Experiments
A.1 Additional evaluation benchmarks.
Our main paper evaluates OpenVLThinker across six widely-used vision-language benchmarks covering mathematical reasoning, general reasoning, and perceptual reliability. These benchmarks are consistent with those used in recent reports on both proprietary (e.g., GPT, Gemini) and open-source (e.g., Qwen-VL, Intern-VL) models. This setup aligns with concurrent works [huang2025visionr1, chen2025sft] that also employ these benchmarks with emphasis on reasoning ability.
To further clarify benchmark coverage, Table 6 provides subset-level EMMA results, showing performance across Math, Chemistry, Physics, and Code categories.
Table 6: Subset performance on EMMA benchmark.
Model EMMA-Math EMMA-Chemistry EMMA-Physics EMMA-Code
Qwen2.5-VL 24.6 21.9 29.5 28.0
VLAA-Thinker 28.1 22.3 28.8 27.3
OpenVLThinker 28.8 22.6 32.7 26.8
In addition, we expanded the evaluation to include two recent benchmarks, MM-Star [chen2024we] and WeMath [qiao2024we]. MM-Star assesses six major LVLM capabilities, including fine-grained perception, mathematics, science & technology, and logical reasoning.
Table 7: Results on newly included benchmarks $M$-Star and WeMath.
Model $M$-Star WeMath
Qwen2.5-VL 53.9 61.9
VLAA-Thinker 55.4 62.4
OpenVLThinker 61.9 64.1
A.2 Computational Cost of the Iterative SFT–RL Loop
[…]
Stage GPU Hours (per GPU) Total (8 GPUs)
SFT 0.06 (3 min 30 s) 0.48
GRPO-Medium 2.01 (2 h 35 s) 16.08
GRPO-Hard 4.57 (4 h 34 min 26 s) 36.56
The SFT stage incurs minimal computational cost due to the small dataset size (3k examples). For RL training on medium-difficulty data, the number of epochs is reduced to maintain efficiency. Although the hard-stage RL incurs the highest cost, the overall compute remains comparable to contemporary RL-based post-training methods.
Importantly, the preceding SFT and medium RL stages accelerate convergence during the final RL stage. Despite the iterative nature introducing additional overhead, the total compute remains practical and resource-efficient for academic-scale training. We plan to include precise GPU-hour estimates and discuss scalability trade-offs in future versions.
A.3 Single-Stage SFT-Only and RL-Only Baselines
To isolate the effect of iteration beyond simply combining SFT and RL, we conducted experiments with single-stage SFT-only and RL-only baselines trained on the same 12K examples used in OpenVLThinker.
  • RL-only: Qwen2.5-VL trained exclusively with GRPO on the full 12K dataset.
  • SFT-only: Qwen2.5-VL trained solely via SFT on iteration-2 trajectories generated by OpenVLThinker. Instances with no correct reasoning within $k=4$ samplings were filtered.
[…]
Table 9: Comparison of single-stage baselines and iterative OpenVLThinker.
Method MathVista EMMA HallusionBench
RL-Only 71.3 24.5 66.8
SFT-Only 71.1 22.3 65.4
OpenVLThinker (Iterative) 71.7 25.2 70.2
[…]
A.4 Impact of Caption Quality on Iterative Training
The quality of caption-based SFT data significantly influences reasoning performance in the iterative training loop. During iteration 1, we constructed the dataset using captions generated by QwQ-32B, filtered by final-answer correctness. Higher-quality captions increased the likelihood of correct reasoning traces, thus expanding the effective training pool.
To analyze this effect, we compared two variants: one using captions from the weaker Qwen2.5-VL-3B model and another using Qwen2.5-VL-7B, both under identical rejection-sampling conditions ($k=4$). The performance evolution across iterations on the MathVista benchmark is shown below.
Table 10: Effect of caption quality across training iterations on MathVista.
Caption Source SFT-Iter1 GRPO-Iter1 SFT-Iter2 GRPO-Iter2 SFT-Iter3 GRPO-Iter3
3B Caption 62.5 65.6 66.1 69.4 69.0 70.2
7B Caption 63.4 66.6 67.5 70.9 69.5 71.7
Better caption quality improves visual grounding and reasoning trace precision in early stages, yielding higher-quality data for subsequent iterations. Consequently, richer initial captions amplify the benefits of the iterative framework, leading to more consistent improvements in reasoning performance.
Appendix B Additional Empirical Study
@@
Does Complex Reasoning Matter for VQA? We additionally investigated whether complex, multi-step reasoning provides significant performance gains over standard (non-R1) reasoning in visual tasks. In this study, we use the ConTextual [wadhawan2024contextual] validation set of 100 VQA examples, aiming to disentangle the roles of image grounding and textual reasoning. As similar to our first distillation process, we separately employ a vision-language model for caption generation and a pure-text model for reasoning. The image description generated by the captioning model is then fed into one of two text-based models: DeepSeek-R1-Distill-14B (an R1-style reasoner) or Qwen2.5-14B-Instruct (a standard instruction-tuned model). This setup allows us to isolate the impact of R1 reasoning from the effects of the underlying vision encoder.
@@
Appendix C Experiment Details
We thank LLaMA-Factory [^6] https://github.com/hiyouga/LLaMA-Factory and EasyR1777https://github.com/hiyouga/EasyR1 for open-sourcing the training framework that we used for SFT and GRPO. In Table 13 and 14, we detail the hyperparameters that we used for SFT, GRPO and inference. We further lay out the prompts we used for generating image captions. Experiments were conducted on GPU clusters to the similar level of NVIDIA H100 80GB GPU. SFT/Distillation requires 30 minutes and RL requires 20 hours for each iteration. In addition, distillation data generation with verification requires about 8 hours.
@@ Appendix C Experiment Details
Appendix D Output Examples
@@
Example of SFT-Iter3 incorrect output and GRPO-Iter3 Correct output to the same question Question: The derivative of f(x) at x=2 is $__$ that at x=5
Choices: (A) larger than (B) equal to (C) smaller than
SFT-Iter3
@@
Okay, let me see. The question is asking about the derivative of the function f(x) = |2x - 3| + 1 at x=2 compared to x=5. First, I need to recall what the derivative represents. The derivative gives the slope of the tangent line at a point on the function. So I need to look at the slopes at those two points. </think> The function ( f(x) = |2x - 3| + 1 ) is an absolute value function, which means it has a V-shape. The vertex of this V-shape occurs where the expression inside the absolute value is zero, i.e., ( 2x - 3 = 0 ), which gives ( x = $\frac{3}{2}$ ). The derivative of ( f(x) ) will be zero at the vertex because the function changes direction there. For ( $x<\frac{3}{2}$ ), the function is decreasing, and for ( $x>\frac{3}{2}$ ), it is increasing. The slopes on either side of the vertex are negative and positive, respectively. At ( x = 2 ) (which is to the right of the vertex), the function is increasing, so the derivative is positive. At ( x = 5 ) (which is further to the right), the function is also increasing, so the derivative is positive but steeper because the function is increasing faster. Since the function is increasing faster at ( x = 5 ) than at ( x = 2 ), the derivative at ( x = 5 ) is larger. <answer>(C)</answer> ================== GRPO-Iter3 <think> Okay, let’s see. The question is asking about the derivative of the function f(x) = |2x - 3| + 1 at x=2 compared to x=5. First, I need to recall what the derivative represents. The derivative at a point gives the slope of the tangent line at that point. So I need to look at the slopes at x=2 and x=5. </think> The function ( f(x) = |2x - 3| + 1 ) is an absolute value function shifted up by 1. The absolute value function ( |2x - 3| ) has a V-shape with a vertex at ( x = $\frac{3}{2}$ ). The derivative of ( |2x - 3| ) is ( 2 ) for ( x $\frac{3}{2}$ ) and ( -2 ) for ( x < $\frac{3}{2}$ ). Adding 1 to the function doesn’t affect the derivative’s magnitude but shifts the y-values up. At ( x = 2 ), which is greater than ( $\frac{3}{2}$ ), the derivative is ( 2 ). At ( x = 5 ), which is also greater than ( $\frac{3}{2}$ ), the derivative is also ( 2 ). Since the derivative is the same at both points, the slopes are equal.
<answer>(B)</answer>
Ground truth: (B)
Example of SFT-Iter3 incorrect output and GRPO-Iter3 Correct output to the same question Question: Find $m\angle 3$.
Choices: (A) 28 (B) 38 (C) 52 (D) 62
SFT-Iter3
@@
Okay, let me see. The problem is asking for the measure of angle 3 in the given quadrilateral. The hint suggests looking at the angles provided, which are 52 degrees and 38 degrees. Hmm, the quadrilateral is a rectangle, so all angles should be 90 degrees. But the hint is confusing me. Let me check the options again. The angles 52 and 38 are given, but they’re not the ones I need to find. The question is about angle 3, which is opposite angle 1. Since opposite angles in a rectangle are equal, angle 3 should be the same as angle 1. But angle 1 is not given directly. Let me see, the hint is saying to look at the angles provided. The 52 and 38 are supplementary angles, which means they add up to 180 degrees. But angle 3 is not supplementary to 52 or 38. It’s the opposite angle, so it should be the same as angle 1. Since angle 1 is not given, I need to use the options. The only angle that could be the same as angle 1 is the one that’s not given directly. The options are A, B, C, D. The only one that’s not supplementary to 52 or 38 is 62. So angle 3 must be 62 degrees. </think> The problem states that the quadrilateral is a rectangle, and all angles in a rectangle are 90 degrees. The hint provided the angles 52° and 38°, which are supplementary angles (52° + 38° = 90°). However, angle 3 is the opposite angle to angle 1, which means they are equal because opposite angles in a rectangle are congruent. Since angle 1 is not given directly, we need to look at the options provided. The only angle that fits the properties of opposite angles in a rectangle is the one that is not supplementary to the given angles. The options are A (28), B (38), C (52), and D (62). Since angle 1 is not given, the only angle that could be the same as angle 3 is the one that is not supplementary to 52° or 38°. The only angle that fits this criterion is 62°, as it is the only one not listed as supplementary to the given angles. <answer>(D)</answer> ================== GRPO-Iter3 <think> Okay, let’s see. The question is asking for the measure of angle 3. The image provided is a rectangle with some angles labeled. The hint suggests looking at the angles given, which are 52 degrees and 38 degrees. Hmm, the rectangle has right angles, so each angle should be 90 degrees. The angles 52 and 38 are supplementary because they add up to 90 degrees (90 - 52 = 38). But angle 3 is not one of those. It’s the angle opposite the 38-degree angle, so it should be the same. Wait, the question is about angle 3, which is the one opposite the 38-degree angle. So it should be 38 degrees too, right?
</think>
The problem states that angle 3 is opposite the 38-degree angle. In a rectangle, opposite angles are equal. Since angle 3 is opposite the 38-degree angle, angle 3 must also be 38 degrees. <answer>(B)</answer> Ground truth: (B)